Telescope Bootstrap Startup: 8 September 2012 UT

Observer: JHill (LBTO)
Telescope Operator: DHuerta (LBTO)
Software Support: CBiddick (Tucson)
AO Support: DMiller (Tucson)
Telescope Support: JUrban (LBTO)
Instrument Support: JMorris (LBTO)

Summary

Heavy Thunderstorms at Sunset. This night has also been lost to weather, but we got lots of other interesting stuff done.

Problems remaining
  • SX Adsec Electronics Shutdown when they get too warm
  • Can't Communicate with AGW2
  • SX M3 doesn't move when we authorize
  • Can't load SX AdSec programs
  • One bad actuator (cable to dist board?) on DX primary)

Problems solved
  • Converted Pointingstar5 to F95 and have it running.
  • Closed dome presets completed
  • M3 swing arms fully operational
  • Fixed Perl scripts used by DOFPIA to talk to LBC

Not Tested
  • Mirror Ventilation
  • Shutter Doors
  • Vent Doors

Details (UT Times)

00:00 Cleaning the control room.

01:00 JHoward and JUrban inspecting both of the M3 swing arms.

01:30 CBiddick and MDelapena are running authorizing tests on 2012G (one of few closed-dome activities that we didn't finish last night.)

01:40 JMorris is removing the MODS1 collimator covers

02:19 Torsten can't ping MOXA in AGW2 reliably.

02:30 Cycling AGW2 power doesn't help.

04:00 We loose connection with SX M2 hexapod again. We move the telescope to horizon to check the AdSec electronics box. The electronics box looks happy - not turned off by thermal switch or loss of communication. Hmmmm? Communication with hexapod returned while we were moving the telescope.

05:00 Cycling power 4 more times on AGW2 didn't help. We did verify that the AGW had power by plugging in an alarm clock at the conveniently provided socket.

05:35 Turning on LBC failed. Reboot lbccontrol

05:46 Successful LBC preset with dome closed

06:15 Perl scripts that DOFPIA uses to talk to LBC are not working.

IDL problems again with SX Adaptive Secondary.

06:30 SX AdSec shuts off again. This time we catch it with the red light on that indicates a thermal cutoff.

07:01 SX Adsec rack has turned back on.

07:15 The problem in the DOFPIA perl scripts was an obsolete URL. I changed lbccontrol.lbt.as.arizona.edu to lbccontrol.

07:53 Doug can't get the program for the SX secondary to load. He's added a note to Issue 4045. This is unrelated to Issue 4188. Runa can look at it in the morning.

08:01 An authorize for LBTI on both sides fails to move the SX tertiary.

08:05 Stopped and restarted OSS to be really sure I've got the patch for V2.27.1 running

09:40 Converted Pointingstars5 to f95, and have it running from /home/LBTO/POINTING/

Details

-- JohnHill - 07 Sep 2012
Topic revision: r4 - 08 Sep 2012, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback